a car going a speed of 44 m/s drops a rock from a
height of 2.2 m How far from the point of the drop will the rock hit the ground?The acceleration of
gravity is 9.8 m/s2
1
Expert's answer
2012-10-31T13:11:13-0400
Let's find the fall time of a rock T:

H = gT²/2 ==> T = √(2H/g) = √(2*2.2[m]/9.8[m/s²]) ≈ 0.45 [s].

So, a rock will drop

D = 0.45[s]*44[m/s] = 19.8 meters

far from the point of the drop.